OpenStudy (anonymous):

A machine is rolling a metal cylinder under pressure. The radius of the cylinder is decreasing at a constant rate of 0.05 in per second and the volume V is 128pi cubic in. At what rate is the length h changing when the radius r is 1.8 in?

OpenStudy (earthcitizen):

\[V=A*h\]

OpenStudy (earthcitizen):

can you state options to the question ?

OpenStudy (mathmate):

\[V=\pi r^2h\] Assuming volume is constant (i.e. material incompressible, differentiate both sides with respect to t (time) \[\frac{dV}{dt} = \pi r(2\frac{dr}{dt}+r \frac{dh}{dt})\] Since V is constant, dV/dt =0, so we get \[2\frac{dr}{dt}+r \frac{dh}{dt} = 0\] Now r and dr/dt are both known, can you solve for dh/dt?

OpenStudy (mathmate):

Sorry, the first equation should read: \[\frac{dV}{dt} = \pi r(2h \frac{dr}{dt}+r \frac{dh}{dt})\] and the second: \[2h \frac{dr}{dt} + r \frac{dh}{dt} = 0\] h can be calculated from V=pi r^2h h=V/(pi r^2)

OpenStudy (earthcitizen):

@mathmate, what methods was used here ?

OpenStudy (mathmate):

It is basically partial differentiation, since both h and r are functions of time. The differentiation is done similar to the differentiation of products, d(uv)=vdu+udv while holding the other variable constant.

OpenStudy (earthcitizen):

the rhs was derived by implicit differentiation ?

OpenStudy (mathmate):

It is slightly different. Implicit differentiation usually involves only two variables, while here we have a third variable which is the independent variable, both x and y are dependent on t. The tree diagram looks like: V - r - t \ h - t That is why we need to differentiate both r and h with respect to t, keeping the other variable constant.

OpenStudy (mathmate):

Perhaps a Wiki article can explain better than I can: http://en.wikipedia.org/wiki/Partial_derivative I apologize for not having used the partial derivative notaion as in: \[2h \frac{\partial r}{\partial t} + r \frac{\partial h}{\partial t} = 0\]

OpenStudy (earthcitizen):

smashing!